How far will a car travel in 30 minutes at 10 miles per second?
Answer by Alan3354(69443)

How far will a car travel in 30 minutes at 10 miles per second?
-----------
30 mins = 1800 seconds
d = r*t = 10*1800
d = 18000 miles
Fast car